The current I in an electrical conductor vary inversely as the resistance R of the conductor. The current is 5 amperes while the resistance is 950 ohms. What is the current when the resistance is 510 ohms?
Answer by Alan3354(69443)

The current I in an electrical conductor vary inversely as the resistance R of the conductor. The current is 5 amperes while the resistance is 950 ohms. What is the current when the resistance is 510 ohms?
-----------
V = I*R = 5*950
V = 4750
--------
4750 = I*510
I = 9.314 amps